# Bridgeland/weak Stability Conditions under Spherical Twist Associated to a Torsion Sheaf

### v1

@@ 5. Bridgeland stability condition under $\text{ST}_{\mathcal{O}_{C}(-1)}$
Let $X$ be a K3 surface, $\nu_{a}\in\text{NS}_{\mathbb{Q}}(X)$ be an arbitrary ample class. Let $\sigma_{V,\nu_{a}}=(Z_{V,\nu_{a}},\mathrm{Coh}^{\nu_{a}})$ be the standard Bridgeland stability condition associated to $\nu_{a}$. In this Section, we study the Bridgeland stability condition $\text{ST}_{\mathcal{O}_{C}(-1)}(\sigma_{V,\nu_{a}})$. A particularly interesting example is as follows. Let $C$ and $\nu$ be as in the previous Section, taking $D=\frac{1}{e}(\nu-C)$, then $\nu$ can be written as $\nu=C+eD$ with $D\cdot C=1$. Assume $D$ is a nef divisor. Let $\nu_{a}=\nu+aD$. If $a>0$, $\nu_{a}$ is ample. If $a<0$, then $\nu_{a}\cdot C<0$, hence $\nu_{a}$ is not nef. Then the construction in this Section gives a description of the Bridgeland stability condition associated to a non-nef divisor. This is our original motivation, so we keep the notation of the ample class to be $\nu_{a}$, but the construction in this Section works for an arbitrary ample class in $\text{NS}_{\mathbb{Q}}(X)$ such that there exists a solution to the central charge formula 4.0.1 in Section 4.
@@
$S^{a}_{k}:\{E\in\mathrm{Coh}^{\nu_{a}}|E\hookrightarrow\mathcal{O}_{C}(j)\text{ for some }j\leq k\}.$
@@
Using standard argument, we have the following simple Lemma. Lemma 5.1. Let $\mathcal{A}$ be an abelian category, and $\mathcal{A}^{\prime}$ a subcategory of $\mathcal{A}$ closed under extension. Let
@@
Let $\mathcal{B}_{\nu_{a},k}$ be the abelian category by tilting $\mathrm{Coh}^{\nu_{a}}$ at $(\mathcal{T}^{1}_{\nu_{a},k},\mathcal{F}^{1}_{\nu_{a},k})$, i.e. $\mathcal{B}_{\nu_{a},k}=\langle\mathcal{F}^{1}_{\nu_{a},k}[1],\mathcal{T}^{1}_{\nu_{a},k}\rangle$ Note that Proposition 10.6 in [6] implies that $B_{\nu_{a},k}$ can be obtained by tilting $\mathrm{Coh}(X)$ at the torsion pair $(\mathcal{T}_{\nu_{a},k},\mathcal{F}_{\nu_{a},k})$, where
@@
$0\to\ker(f)\to\mathcal{O}_{C}(k)[1]\to\text{im}(f)\to 0$
where $\text{im}(f)\in\mathcal{F}_{\nu_{a},k}$. Hence $\text{im}(f)[-1]$ has no zero dimensional subsheaf. Taking l.e.s. of cohomology, if $H^{-1}(\ker(f))\neq 0$, we must have $H^{-1}(\ker(f))\simeq\mathcal{O}_{C}(k)$. Then $\text{im}(f)[-1]\simeq H^{0}(\ker(f))\in\mathcal{T}_{\nu_{a},k}$, which is a contradiction. So we must have $H^{-1}(\ker(f))=0$, and we have a s.e.s. in $\mathrm{Coh}(X)$
(5.3.1) $0\to\mathcal{O}_{C}(k)\to\text{im}(f)[-1]\to\ker(f)\to 0$
@@
where $J_{i}/J_{i-1}$’s are subobject of $\mathcal{O}_{C}(k)$ in $\mathrm{Coh}^{\nu_{a}}$. Taking the smallest $i$ such that $\textnormal{Hom}(\mathcal{O}_{C}(k),J_{i})\neq 0$. Then $\mathcal{O}_{C}(k)\simeq J_{i}/J_{i-1}$, and $\mathcal{O}_{C}(k)$ is a direct summand of $J_{i}$, then 5.3.1 implies that $\ker(f)\in\mathcal{C}^{a}_{k}$, contradicting $\textnormal{Hom}(\ker(f),\mathcal{C}^{a}_{k})=0$. Hence $\ker(f)=0$ and $f$ is injective.
@@
Let $\mathcal{B}^{o}_{\nu_{a},k}:=\langle\mathcal{F}^{o}_{k},\mathcal{T}^{o}_{k}[-1]\rangle$ be the heart obtained by tiling $\mathcal{B}_{\nu_{a},k}$ at the torsion pair in Corollary 5. Proposition 5.5. Let $f:\mathcal{O}_{C}(k)\to E$ be a nonzero morphism in $\mathcal{B}^{o}_{\nu_{a},k}$, then $f$ is injective.
@@
$0\to\ker(f)\to\mathcal{O}_{C}(k)\to\text{im}(f)\to 0$
Here $\ker(f)$ and $\text{im}(f)$ fit into s.e.s’ in $\mathcal{B}^{o}_{\nu_{a},k}$:
@@
$0\to I_{1}\to\text{im}(f)\to I_{2}\to 0,$
where $K_{1}$ and $I_{1}$ are in $\mathcal{F}^{o}_{k}$, $K_{2}$ and $I_{2}$ are in $\mathcal{T}^{o}_{k}[-1]$. If $K_{1}\neq 0$, since $K_{1}\in\mathcal{T}_{\nu_{a},k}$, we have $\textnormal{Hom}(K_{1},\mathcal{O}_{C}(k))=0$. Then $K_{1}$ is a subsheaf of $H^{-1}(\text{im}(f))$. Since $H^{0}(\text{im}(f))$ is the zero sheaf or supported in dimension 00, we have $I_{2}=0$ hence $\text{im}(f)\simeq I_{1}$. Then $H^{-1}(\text{im}(f))\in\mathcal{F}_{\nu_{a},k}$, which implies that $K_{1}\in\mathcal{F}_{\nu_{a},k}$. This contradicts $K_{1}\in\mathcal{T}_{\nu_{a},k}$, hence $K_{1}=0$. But this will force $\ker(f)\in\langle\mathcal{O}_{C}(k)\rangle$, which also leads to a contradiction.
@@
[…]
@@
$\begin{split}\mathcal{T}^{s}_{\nu_{a},-1}&=\{E\in\text{ST}(\mathcal{B}_{\nu_{a},-1})|\forall E\twoheadrightarrow F\in\text{ST}(\mathcal{B}_{\nu_{a},-1}),\Im Z_{V,\zeta}(F)\leq 0\}\\ \mathcal{F}^{s}_{\nu_{a},-1}&=(\mathcal{T}^{s}_{\nu_{a},-1})^{\perp}.\end{split}$
@@
and $\mathcal{A}_{{\nu_{a}},-1}=(\text{ST}(\mathcal{F}^{o}_{-1}),\text{ST}(\mathcal{T}^{o}_{-1})[-1])$, we naturally have
$\begin{split}\mathcal{F}^{s}_{-1}&=\text{ST}(\mathcal{T}^{o}_{-1})[-1]=\langle\mathcal{O}_{C}(-1)[-1]\rangle\\ \mathcal{T}^{s}_{-1}&=\{T\in\mathcal{A}_{{\nu_{a}},-1}|\textnormal{Hom}(T,\mathcal{F}^{s}_{-1})=0\}.\end{split}$
@@
(5.8.1) $\mathcal{T}_{2}=\{E\in\mathcal{B}_{\nu_{a},-1}|\forall E\twoheadrightarrow F\in\mathcal{B}_{\nu_{a},-1},\Im Z_{V,\nu_{a}}(F)\leq 0.\}$
If $E\in\mathcal{T}_{2}$, for any $E\twoheadrightarrow F$ in $\mathcal{B}_{\nu_{a},-1}$, $F$ is also in $\mathcal{T}_{2}$, i.e. $F\in\mathcal{F}^{1}_{\nu_{a},-1}[1]$. Then $\Im Z_{V,\nu_{a}}(F)\leq 0$. Conversely, if $E\notin\mathcal{T}_{2}$, then $E$ fits into a s.e.s. in $\mathcal{B}_{\nu_{a},-1}$:
@@
where $E_{1}\in\mathcal{F}^{1}_{\nu_{a},-1}[1]$, $E_{2}\in\mathcal{T}^{1}_{\nu_{a},-1}$ and $E_{2}\neq 0$. But $\Im Z_{V,\nu_{a}}(E_{2})>0$, hence $E$ is not an element of the RHS of 5.8.1.
Define $g\in\tilde{\text{GL}}^{+}(2,\mathbb{R})=(g_{0},f)$, where $f$ is determined by the property that for any $0<\phi\leq 1$, $0<f(\phi)\leq 1$. Corollary 5.9.Let $\mathcal{A}^{\nu_{a}}$ be the abelian category obtain by tilting $\mathrm{Coh}^{\nu_{a}}$ at the sequence of torsion pairs in the first row of 5.7.1, followed by 5.7.2. Then $(Z_{V,\zeta},\mathcal{A}^{\nu_{a}})$ defines a Bridgeland stability conditions and
@@
Remark 5.10.  Let $C$, $\nu$ and $D$ be as in the first paragraph of this Section. Take $\nu_{a}=\nu+aD$ for some $a\in\mathbb{Q}_{>0}$. Assuming $D$ is nef, then $\nu_{a}$ is ample. Then the Bridgeland stability condition $\text{ST}_{\mathcal{O}_{C}(-1)}(\sigma_{V,\nu_{a}})$ is of the form $\sigma_{V,\nu_{b}}=(Z_{V,\nu_{b}},\text{ST}_{\mathcal{O}_{C}(-1)}(\mathrm{Coh}^{\nu_{a}}))$. From the computation in Section 4.1, we have $\nu_{b}=\nu+bD$ where $b=-\frac{a}{a+1}$. Then Proposition 5 gives an explicit construction of Bridgeland stability condition associated to a non-nef divisor.
Finally, we show that the Bridgeland stability condition $\text{ST}_{\mathcal{O}_{C}(-1)}(\sigma_{V,\nu_{a}})$ is not in the closure of the geometric chamber of the stability manifold. We first compute the image of skyscraper sheaves under spherical twist. Proposition 5.11. Let $p$ be a point in $X$.
@@
(i) follows directly from the definition of spherical twist.
For (ii), GRR implies that $H^{0}(\mathcal{O}_{C}(1))-H^{1}(\mathcal{O}_{C}(1))=2$. Since $H^{0}(\mathcal{O}_{C}(1))$ is two dimensional, we must have $H^{1}(\mathcal{O}_{C}(1))=0$. By assumption $e=-C^{2}=2$, which implies that $H^{0}(\mathcal{O}(-C)^{*}\otimes\mathcal{O}_{C}(1))=0$. Resolving $\mathcal{O}_{C}$ using $\mathcal{O}_{X}(-C)\to\mathcal{O}_{X}$, then local to global spectral sequence implies that $\textnormal{Ext}^{1}(\mathcal{O}_{C},\mathcal{O}_{C}(1))=0$. Note that $Li^{*}i_{*}\mathcal{O}_{C}\simeq(Li^{*}i_{*}\mathcal{O}_{C}(t))\otimes\mathcal{O}_{C}(-t)$, using adjoint functors, for any $j$ we have
$\textnormal{Ext}^{j}(\mathcal{O}_{C}(t),\mathcal{O}_{C}(t+1))=\textnormal{Ext}^{j}(\mathcal{O}_{C},\mathcal{O}_{C}(1)).$
From the above discussion we have $\text{ST}_{\mathcal{O}_{C}(t)}(\mathcal{O}_{C}(t+1))=\mathcal{O}_{C}(t-1)[1]$. Also from the definition of spherical twist, we have $\text{ST}_{\mathcal{O}_{C}(t)}(\mathcal{O}_{C}(t))=\mathcal{O}_{C}(t)[-1]$. Applying $\text{ST}_{\mathcal{O}_{C}(t)}^{-1}$ to the exact triangle
$0\to\mathcal{O}_{C}(t-1)\to\mathcal{O}_{C}(t)\to\mathcal{O}_{p}\to 0$
we obtain the following triangle
$\mathcal{O}_{C}(t+1)[-1]\to\mathcal{O}_{C}(t)[1]\to E^{\bullet}\xrightarrow{+1}.$
Taking l.e.s. of cohomology, we obtain the cohomology of $E^{\bullet}$.
A similar computation for the s.e.s
$0\to\mathcal{O}_{C}(t)\to\mathcal{O}_{C}(t+1)\to\mathcal{O}_{p}\to 0$
gives the cohomology for $\text{ST}_{\mathcal{O}_{C}(t)}(\mathcal{O}_{p})$.

### v2

@@ 5. Bridgeland stability condition under $\text{ST}_{\mathcal{O}_{C}(-1)}$
Let $X$ be a K3 surface, $\nu_{a}\in\text{NS}_{\mathbb{Q}}(X)$ be an arbitrary ample class. Let $\sigma_{V,\nu_{a}}=(Z_{V,\nu_{a}},\mathrm{Coh}^{\nu_{a}})$ be the standard Bridgeland stability condition associated to $\nu_{a}$. In this Section, we study the Bridgeland stability condition $\text{ST}_{\mathcal{O}_{C}(-1)}(\sigma_{V,\nu_{a}})$. A particularly interesting example is as follows. Let $C$ and $\nu$ be as in the previous Section, taking $D=\frac{1}{e}(\nu-C)$, then $\nu$ can be written as $\nu=C+eD$ with $D\cdot C=1$. Assume $D$ is a nef divisor. Let $\nu_{a}=\nu+aD$. If $a>0$, $\nu_{a}$ is ample. If $a<0$, then $\nu_{a}\cdot C<0$, hence $\nu_{a}$ is not nef. Then the construction in this Section in particular constructs a Bridgeland stability condition associated to a non-nef divisor. This is our original motivation, so we keep the notation of the ample class to be $\nu_{a}$, but the construction in this Section works for an arbitrary ample class in $\text{NS}_{\mathbb{Q}}(X)$ such that there exists a solution to the central charge formula 4.0.1 in Section 4.
@@
$S^{a}_{k}:\{E\in\mathrm{Coh}^{\nu_{a}}|E\hookrightarrow\mathcal{O}_{C}(k)\}.$
@@
Using standard argument, we have the following simple Lemma.
Lemma 5.1. Let $\mathcal{A}$ be an abelian category, and $\mathcal{A}^{\prime}$ a subcategory of $\mathcal{A}$ closed under extension. Let
@@
Let $\mathcal{B}_{\nu_{a},k}$ be the abelian category by tilting $\mathrm{Coh}^{\nu_{a}}$ at $(\mathcal{T}^{1}_{\nu_{a},k},\mathcal{F}^{1}_{\nu_{a},k})$, i.e. $\mathcal{B}_{\nu_{a},k}=\langle\mathcal{F}^{1}_{\nu_{a},k}[1],\mathcal{T}^{1}_{\nu_{a},k}\rangle$ . Note that Proposition 10.6 in [7] implies that $B_{\nu_{a},k}$ can be obtained by tilting $\mathrm{Coh}(X)$ at the torsion pair $(\mathcal{T}_{\nu_{a},k},\mathcal{F}_{\nu_{a},k})$, where
@@
$0\to\ker(f)\to\mathcal{O}_{C}(k)[1]\to\text{im}(f)\to 0.$
where $\text{im}(f)[-1]\in\mathcal{F}_{\nu_{a},k}$. Hence $\text{im}(f)[-1]$ has no zero dimensional subsheaf. Taking l.e.s. of cohomology, if $H^{-1}(\ker(f))\neq 0$, we must have $H^{-1}(\ker(f))\simeq\mathcal{O}_{C}(k)$. Then $\text{im}(f)[-1]\simeq H^{0}(\ker(f))\in\mathcal{T}_{\nu_{a},k}$, which is a contradiction. So we must have $H^{-1}(\ker(f))=0$, and we have a s.e.s. in $\mathrm{Coh}(X)$
(5.3.1) $0\to\mathcal{O}_{C}(k)\to\text{im}(f)[-1]\to\ker(f)\to 0.$
@@
where $J_{i}/J_{i-1}$’s are subobject of $\mathcal{O}_{C}(k)$ in $\mathrm{Coh}^{\nu_{a}}$. Taking the smallest $i$ such that $\textnormal{Hom}(\mathcal{O}_{C}(k),J_{i})\neq 0$. Then $\mathcal{O}_{C}(k)\simeq J_{i}/J_{i-1}$, and $\mathcal{O}_{C}(k)$ is a direct summand of $J_{i}$. Then 5.3.1 implies that $\ker(f)\in\mathcal{C}^{a}_{k}$, contradicting $\textnormal{Hom}(\ker(f),\mathcal{C}^{a}_{k})=0$. Hence $\ker(f)=0$ and $f$ is injective.
@@
Let $\mathcal{B}^{o}_{\nu_{a},k}:=\langle\mathcal{F}^{o}_{k},\mathcal{T}^{o}_{k}[-1]\rangle$ be the heart obtained by tiling $\mathcal{B}_{\nu_{a},k}$ at the torsion pair in Corollary 5.
Proposition 5.5. Let $f:\mathcal{O}_{C}(k)\to E$ be a nonzero morphism in $\mathcal{B}^{o}_{\nu_{a},k}$, then $f$ is injective.
@@
$0\to\ker(f)\to\mathcal{O}_{C}(k)\to\text{im}(f)\to 0.$
Here $\ker(f)$ fits into s.e.s’ in $\mathcal{B}^{o}_{\nu_{a},k}$:
@@
where $K_{1}$ is in $\mathcal{F}^{o}_{k}$, $K_{2}$ is in $\mathcal{T}^{o}_{k}[-1]$. By construction $K_{1}\in\mathcal{T}_{\nu_{a},k}$, hence $\textnormal{Hom}(K_{1},\mathcal{O}_{C}(k))=0$. This forces $K_{1}=0$ and $\ker(f)\in\langle\mathcal{O}_{C}(k)\rangle$, which leads to a contradiction.
@@
(5.7.1) $\mathrm{Coh}^{\nu_{a}}$$\mathcal{B}_{\nu_{a},-1}$$\mathcal{B}^{o}_{\nu_{a},-1}$$\text{ST}_{\mathcal{O}_{C}(-1)}(\mathrm{Coh}^{\nu_{a}})$$\text{ST}_{\mathcal{O}_{C}(-1)}(\mathcal{B}_{\nu_{a},-1})$$\mathcal{A}_{\nu_{a},-1}=\text{ST}_{\mathcal{O}_{C}(-1)}(\mathcal{B}^{o}_{\nu_{a},-1})$$\scriptstyle{(\mathcal{T}^{1}_{\nu_{a},-1},\mathcal{F}^{1}_{\nu_{a},-1})}$$\scriptstyle{(\mathcal{T}^{o}_{-1},\mathcal{F}^{o}_{-1})}$$\scriptstyle{(\mathcal{T}^{c}_{-1},\mathcal{F}^{c}_{-1})}$$\scriptstyle{{\text{ST}_{\mathcal{O}_{C}(-1)}(\mathcal{T}^{1}_{\nu_{a},-1},\mathcal{F}^{1}_{\nu_{a},-1})}}$$\scriptstyle{\text{ST}_{\mathcal{O}_{C}(-1)}(\mathcal{T}^{o}_{-1},\mathcal{F}^{o}_{-1})}$
@@
$\begin{split}\mathcal{T}^{s}_{\nu_{a},-1}&=\{E\in\text{ST}(\mathcal{B}_{\nu_{a},-1})|\forall E\twoheadrightarrow F\in\text{ST}(\mathcal{B}_{\nu_{a},-1}),\Im Z_{V,\zeta}(F)<0\}\\ \mathcal{F}^{s}_{\nu_{a},-1}&=(\mathcal{T}^{s}_{\nu_{a},-1})^{\perp}.\end{split}$
@@
and $\mathcal{A}_{{\nu_{a}},-1}=\langle\text{ST}(\mathcal{F}^{o}_{-1}),\text{ST}(\mathcal{T}^{o}_{-1})[-1]\rangle$, we naturally have
$\begin{split}\mathcal{F}^{s}_{-1}&=\text{ST}(\mathcal{T}^{o}_{-1})[-1]=\langle\mathcal{O}_{C}(-1)[-1]\rangle,\\ \mathcal{T}^{s}_{-1}&=\{T\in\mathcal{A}_{{\nu_{a}},-1}|\textnormal{Hom}(T,\mathcal{F}^{s}_{-1})=0\}.\end{split}$
@@
(5.8.1) $\mathcal{T}_{2}=\{E\in\mathcal{B}_{\nu_{a},-1}|\forall E\twoheadrightarrow F\in\mathcal{B}_{\nu_{a},-1},\Im Z_{V,\nu_{a}}(F)<0.\}$
If $E\in\mathcal{T}_{2}$, for any $E\twoheadrightarrow F$ in $\mathcal{B}_{\nu_{a},-1}$, $F$ is also in $\mathcal{T}_{2}$, i.e. $F\in\mathcal{F}^{1}_{\nu_{a},-1}[1]$. Then $\Im Z_{V,\nu_{a}}(F)<0$. Conversely, if $E\notin\mathcal{T}_{2}$, then $E$ fits into a s.e.s. in $\mathcal{B}_{\nu_{a},-1}$:
@@
where $E_{1}\in\mathcal{F}^{1}_{\nu_{a},-1}[1]$, $E_{2}\in\mathcal{T}^{1}_{\nu_{a},-1}$ and $E_{2}\neq 0$. But $\Im Z_{V,\nu_{a}}(E_{2})\geq 0$, hence $E$ is not an element of the RHS of 5.8.1.
Define $g\in\tilde{\text{GL}}^{+}(2,\mathbb{R})=(g_{0},f)$, where $f$ is determined by the property that for any $0<\phi\leq 1$, $0<f(\phi)\leq 1$.
Corollary 5.9.Let $\mathcal{A}^{\nu_{a}}$ be the abelian category obtain by tilting $\mathrm{Coh}^{\nu_{a}}$ at the sequence of torsion pairs in the first row of 5.7.1, followed by 5.7.2. Then $(Z_{V,\zeta},\mathcal{A}^{\nu_{a}})$ defines a Bridgeland stability conditions and
@@
Remark 5.10.  Let $C$, $\nu$ and $D$ be as in the first paragraph of this Section. Take $\nu_{a}=\nu+aD$ for some $a\in\mathbb{Q}_{>0}$. Assuming $D$ is nef, then $\nu_{a}$ is ample. Then the Bridgeland stability condition $\text{ST}_{\mathcal{O}_{C}(-1)}(\sigma_{V,\nu_{a}})$ is of the form $\sigma_{V,\nu_{b}}=(Z_{V,\nu_{b}},\text{ST}_{\mathcal{O}_{C}(-1)}(\mathrm{Coh}^{\nu_{a}}))$. From the computation in Section 4.1, we have $\nu_{b}=\nu+bD$ where $b=-\frac{a}{a+1}$. Then Proposition 5 gives an explicit construction of a Bridgeland stability condition associated to a non-nef divisor $\nu_{b}$. This suggests that the Bridgeland stability condition $\text{ST}_{\mathcal{O}_{C}(-1)}(\sigma_{V,\nu_{a}})$ may not be in the closure of the geometric chamber of the stability manifold. We verify this by looking at the stability of skyscraper sheaves. The following computation is well known and follows directly from GRR computations.
Proposition 5.11. Let $p$ be a point in $X$.
@@
We finish this Section by an observation that the functor $\text{ST}_{\mathcal{O}_{C}(-1)}$ is equipped with a moduli interpretation. Recall in [17], the author constructed a moduli space $\mathcal{M}_{\tau}([\mathcal{O}_{x}])$ of $\tau$-stable objects of class $\mathcal{O}_{x}]$, and shows that it is isomorphic to $Y:=X\sqcup_{C}\mathbb{P}^{1}$. Since $Y\setminus\mathbb{P}^{1}\simeq X\setminus C$, hence $Y$ is smooth at any point q such that $q\notin\mathbb{P}^{1}$. For a point $p\in\mathbb{P}^{1}$, Lemma 7.10 in [17] implies that $T_{p}Y\simeq\textnormal{Ext}^{1}(E_{p},E_{p})\simeq\mathbb{C}^{2}$. Here $E_{p}$ is the object parametrized by the point $p$. This implies that $Y$ is smooth and hence $X\simeq Y$. Let $\Phi:D^{b}(Y)\to D^{b}(X)$ be the Fourier-Mukai transform with kernel given by the universal object $\mathcal{U}$ on $Y\times X$.
Proposition 5.12.We have an equivalence of functors $\Phi=\text{ST}_{\mathcal{O}_{C}(-1)}^{-1}$.
@@
By Proposition 7.9 in [17], we have
$\mathcal{U}[1]\simeq\text{Cone}(\mathcal{O}_{\Delta_{X}}\to\mathcal{O}_{C\times C}(-1,-1)[1]).$
Proposition 1.1.6 in [10] implies that $\mathcal{E}xt^{i}(\mathcal{O}_{\Delta_{X}},\mathcal{O}_{X\times X})=0$ for $i=0,1$, and $\mathcal{E}xt^{3}(\mathcal{O}_{\Delta_{X}},\mathcal{O}_{X\times X})$ is supported in dimension 0. Since $\textnormal{Ext}^{3}(\mathcal{O}_{\Delta_{X}},\mathcal{O}_{X\times X})=0$, local to global spectral sequence implies that $\mathcal{E}xt^{3}(\mathcal{O}_{\Delta_{X}},\mathcal{O}_{X\times X})=0$. The proof of 1.1.8 in [10] also implies that $\mathcal{O}_{\Delta X}^{\vee}\simeq\mathcal{E}xt^{2}(\mathcal{O}_{\Delta_{X}},\mathcal{O}_{X\times X})$ is pure. Then the Chern character forces $\mathcal{O}_{\Delta_{X}}^{\vee}\simeq\mathcal{O}_{\Delta_{X}}[-2]$. Similar argument shows that $\mathcal{O}_{C}(-1)^{\vee}\simeq\mathcal{O}_{C}(-1)[-1]$, hence we have $\mathcal{O}_{C\times C}(-1,-1)^{\vee}\simeq\mathcal{O}_{C\times C}(-1,-1)[-2]$. Hence, we obtain
$\begin{split}\mathcal{U}[-2]&\simeq\text{Cone}(\mathcal{O}_{\Delta X}^{\vee}\to\mathcal{O}_{C\times C}(-1,-1)^{\vee}[1])[-1]\\ &\simeq\text{Cone}(\mathcal{O}_{C\times C}(-1,-1)[-1]\to\mathcal{O}_{\Delta_{X}})^{\vee}.\end{split}$
Recall in Definition 2.3, $\text{ST}_{\mathcal{O}_{C}(-1)}$ is a Fourier-Mukai transform with kernel equals to
$\text{Cone}(\mathcal{O}_{C\times C}(-1,-1)[-1]\to\mathcal{O}_{\Delta X}),$
hence $\mathcal{U}$ is the kernel of $\text{ST}_{\mathcal{O}_{C}(-1)}^{-1}$. Since both functors are Fourier-Mukai transforms, there is an equivalence of functors $\Phi=\text{ST}_{\mathcal{O}_{C}(-1)}^{-1}$. In particular this also implies that $\Phi$ is an equivalence.